Write an equation or formula that represents the general relationship indicated. Then use the given information to find the unknown value.
If y is directly proportional to x^2 and y = 2 when x = 3, what is y when x = 18?

Answer by Cromlix(4381)

y = kx^2
k = constant of variation
To find k:
Use values given.
y = kx^2
2 = k3^2
2 = k9
k = 2/9
y = 2/9x^2
y = 2/9 * 18^2
y = 72
Hope this helps.
:-)